\begin{thebibliography}{}



\bibitem{Barbour2011}
A. D. Barbour and G. Reinert,
The shortest distance in random multi-type intersection graphs,
Random Structures and Algorithms
39 (2011),
179--209.



\bibitem{behrisch2007}
M.~Behrisch, Component evolution in random intersection graphs, The
Electronical Journal of Combinatorics 14 (2007), $\#$R17.

\bibitem{Blackburn2009}
S. Blackburn and S. Gerke,
Connectivity of the uniform random intersection graph,
Discrete Mathematics
309 (2009),
5130-5140.


\bibitem{Bloznelis2008}
M. Bloznelis,
Degree distribution of a typical vertex in a general random intersection graph,
Lithuanian Mathematical Journal
 48 (2008), 38--45.


\bibitem{Bloznelis2010}
M. Bloznelis,
A random intersection digraph: Indegree and outdegree distributions,
Discrete Mathematics
 310 (2010), 2560--2566.

\bibitem{Bloznelis2010+}
M. Bloznelis,  
The largest component in an inhomogeneous random intersection graph with clustering,
The Electronic Journal of Combinatorics 17(1) (2010), R110. 

 
\bibitem{Bloznelis2011+}
M. Bloznelis,
Degree and clustering coefficient in sparse random intersection graphs, to appear in The Annals of Applied Probability.


\bibitem{BloznelisKurauskas2012}
M. Bloznelis, V. Kurauskas,
Clustering function: a measure of social influence,  http://arxiv.org/abs/1207.4941

\bibitem{Bradonjic2010}
M. Bradonjic, A. Hagberg, N. W. Hengartner, A. G. Percus, Component Evolution in General Random Intersection Graphs, The 7th Workshop on Algorithms and Models for the Web Graph, WAW2010. 
Lecture Notes in Computer Science (Springer-Verlag, Berlin, 2010), Vol. 6516, pp. 36-49. 

\bibitem{Britton2008}
T. Britton, M. Deijfen, M. Lindholm, and N. A. Lageras, 
Epidemics on random graphs with tunable clustering.  J. Appl. Prob. 45  (2008), 743--756.

\bibitem{Deijfen}
    { M. Deijfen and W. Kets},
    {Random intersection graphs with tunable degree distribution and clustering},
Probab. Engrg. Inform. Sci. 23 (2009), 661--674.



\bibitem{eschenauer2002}
   L. Eschenauer  and V. D. Gligor,
   A key-management scheme for distributed  sensor networks,
   in: Proceedings of the $9$th ACM Conference on Computer and Communications
   Security
   (2002),
  41--47.



\bibitem{godehardt2001}
   E. Godehardt and J. Jaworski,
   Two models of random intersection graphs and their applications,
   {\it Electronic Notes in Discrete Mathematics}~{\bf 10} (2001), 129--132.



\bibitem
{JKS}
J. Jaworski, M. Karo\'nski, and D. Stark,
   The degree of a typical vertex in generalized random intersection  graph models,
   Discrete Mathematics
   306 (2006), 2152--2165.

\bibitem
{JaworskiStark2008}
J. Jaworski and D. Stark,
   The vertex degree distribution of passive random intersection graph models,
 {\it Combinatorics, Probability and Computing}~{\bf 17} (2008), 549--558.



   \bibitem{karonski1999}
   M. Karo\'nski, E. R.  Scheinerman, and K. B. Singer-Cohen,
   On random intersection graphs: The subgraph problem,
   Combinatorics, Probability and Computing
   8 (1999),
   131--159.







\bibitem{Spirakis2004}
S. Nikoletseas, C. Raptopoulos, and P. G. Spirakis, The existence and efficient construction
of large independent sets in general random intersection graphs, in ICALP (2004), J. Daz, J. Karhumki,
A. Lepist, and D. Sannella, Eds., vol. 3142 of Lecture Notes in Computer Science, Springer, pp. 1029–1040.





\bibitem{Rybarczyk2011}
 K. Rybarczyk,
Diameter, connectivity, and phase transition of the uniform random intersection graph,
Discrete Mathematics  311 (2011), 1998--2019.



\bibitem{Rybarczyk-degree2011}
K. Rybarczyk,
The degree distribution in random intersection graphs,
 in:
   Studies in Classification, Data Analysis and Knowledge Organization,
   Springer,
   Berlin--Heidelberg--New York,
   (2012), 291--299.




\bibitem{Shang2010}
Y. Shang, Degree distributions in general random intersection graphs, The
Electronical Journal of Combinatorics 17 (2010), $\#$R23.

\bibitem{Singer1995}
K. B. Singer-Cohen, Random intersection graphs, PhD thesis, Department of Mathematical
Sciences, The Johns Hopkins University, 1995


\bibitem{stark2004}
D. Stark,
The vertex degree distribution of random intersection graphs,
Random Structures and Algorithms
24 (2004),
249--258.

\bibitem{Steele}
J. M. Steele,
Le Cam's inequality and Poisson approximations,
The American Mathematical Monthly 
101 (1994), 48--54. 




\end{thebibliography}

\documentclass[11pt,a4paper]{article}
\usepackage{amsmath,amssymb,amsfonts,amsthm}
\usepackage[pdftex]{graphicx, color}

\newcommand{\E}{{\bf{E}}}
\newcommand{\PP}{{\bf{P}}}

\newtheorem{tm}{Theorem}

\newtheorem{lem}{Lemma}

\begin{document}

\bibliographystyle{plain}
\parindent=0pt
\centerline{\LARGE \bfseries Degree distribution of an inhomogeneous}
\centerline{\LARGE \bfseries  random intersection graph}

\par\vskip 3.5em

\centerline{Mindaugas  Bloznelis, Julius Damarackas}

\vglue2truecm

\centerline{Vilnius University, Faculty of mathematics and informatics,} 
\centerline{Naugarduko 24, Vilnius,
03225, Lithuania}

%\footnotetext[*]{Faculty of Mathematics and Informatics, Vilnius University, 03225 Vilnius, Lithuania}
%\centerline{ Mindaugas Bloznelis\footnote{Research supported by the Research Council of Lithuania Grant MIP-053}}

\bigskip

% Faculty of Mathematics and Informatics, Vilnius University,

% Naugarduko 24, LT-03225 Vilnius, Lithuania

%E-mail:   mindaugas.bloznelis$@$mif.vu.lt

%http://www.mif.vu.lt/$\sim$bloznelis
%\par\vskip 3.5em

\begin{abstract}
We show the asymptotic degree distribution of the typical vertex of a sparse inhomogeneous random intersection graph. 

\par
\end{abstract}

\smallskip
{\bfseries key words:}  degree distribution, random graph, random intersection graph, power law
\par\vskip 2.5em

2000 Mathematics Subject Classifications:   05C80,  05C07,  05C82

\par\vskip 2.5em

\section{Introduction}

Let $X_1,\dots, X_m, Y_1,\dots, Y_n$ be independent non-negative random variables such that
each $X_i$ has the  probability distribution $P_1$ and each $Y_j$ has the probability distribution
$P_2$. Given realized values $X=\{X_i\}_{i=1}^m$ and $Y=\{Y_j\}_{j=1}^n$
we define the random bipartite graph $H_{X,Y}$ with the bipartition $V=\{v_1,\dots, v_n\}$, $W=\{w_1,\dots, w_m\}$, 
where  edges $\{w_i,v_j\}$ are inserted 
with probabilities
$p_{ij}=\min\{1, X_iY_j(nm)^{-1/2}\}$ independently 
for each $\{i,j\}\in [m]\times [n]$.
The 
 {\it inhomogeneous} random intersection graph  $G(P_1,P_2, n,m)$ defines the adjacency relation on the  vertex 
set $V$: vertices $v',v''\in V$ are declared adjacent (denoted $v'\sim v''$) whenever
$v'$ and $v''$ have a common neighbour in $H_{X,Y}$. 

 The degree distribution of the typical vertex of the random  graph  $G(P_1,P_2, n,m)$ has been first considered  by
Shang \cite{Shang2010}.
The proof of the main result of  \cite{Shang2010} contains  gaps and the  result is incorrect in the regime where
 $m/n\to\beta\in (0,+\infty)$ as $m,n\to+\infty$. 
 We remark that this regime is of particular importance, because it leads to inhomogeneous graphs
 with the clustering property: the clustering coefficient $\PP(v_1\sim v_2|v_1\sim v_3, v_2\sim v_3)$ is bounded away from zero
 provided that
  $\E X_1^3<\infty$ and $\E Y_1^2<\infty$, see \cite{BloznelisKurauskas2012}. 
The aim of the present paper is to show the asymptotic degree distribution in the case where $m/n\to\beta$ for 
some $\beta\in (0,+\infty)$.

 We consider a sequence of graphs $\{G_n=G(P_1,P_2, n,m)\}$, where $m=m_n\to +\infty$ as $n\to \infty$, and
where $P_1, P_2$ do not depend on $n$.
We denote $a_i=\E X_1^i$, $b_i=\E Y_1^i$.
 By $d(v_j)=d_{G_n}(v_j)$ we denote the degree of a vertex $v_j$ in $G_n$ (the number of vertices adjacent to $v_j$ in $G_n$). We remark that for every $n$ the random variables
 $d_{G_n}(v_1),\dots, d_{G_n}(v_n)$ are identically distributed. In Theorem \ref{T1} below we show the asymptotic distribution of $d(v_1)$.

\begin{tm}\label{T1} Let $m,n\to\infty$. 

(i) Assume that $m/n\to 0$. Suppose that  that $\E X_1<\infty$. Then $\PP(d(v_1)=0)=1-o(1)$.

(ii) Assume that $m/n\to \beta$ for some $\beta\in (0,+\infty)$.
 Suppose  that $\E X_1^2<\infty$ and $\E Y_1<\infty$. Then  $d(v_1)$ converges in distribution to the random variable
\begin{equation}\label{d*1}
d_*=\sum_{j=1}^{\Lambda_1}\tau_j, 
\end{equation}
where $\tau_1,\tau_2,\dots$ are independent 
and identically distributed random variables independent of the random variable $\Lambda_1$.
They are distributed as follows. For $r=0,1,2,\dots$, we have
\begin{equation}\label{d*1++}
\PP(\tau_1=r)=\frac{r+1}{\E\Lambda_2}\PP(\Lambda_2=r+1)
\qquad
{\text{and}}
\qquad
\PP(\Lambda_i=r)=\E \,e^{-\lambda_i}\frac{\lambda_i^r}{r!},
\qquad
i=1,2.
\end{equation}
Here $\lambda_1=Y_1a_1\beta^{1/2}$ and $\lambda_2=X_1b_1\beta^{-1/2}$.

(iii) Assume that $m/n\to+\infty$.
 Suppose that  $\E X_1^2<\infty$ and $\E Y_1<\infty$. 
Then  $d(v_1)$ converges in distribution to a random variable
$\Lambda_3$ having the probability distribution
\begin{equation}\label{d*2}
\PP(\Lambda_3=r)=\E e^{-\lambda_3}\frac{\lambda_3^r}{r!},
\qquad
r=0,1,\dots.
\end{equation}
Here $\lambda_3=Y_1a_2b_1$.
\end{tm}

 {\it Remark 1.} The probability distributions $P_{\Lambda_i}$ of $\Lambda_i$, $i=1,2,3$, are Poisson mixtures. One way to
sample 
from the  distribution $P_{\Lambda_i}$  is to 
generate random variable  $\lambda_i$ and then, given $\lambda_i$, 
to generate Poisson random variable 
with the parameter $\lambda_i$. The realized value of the Poisson random variable 
has the  distribution $P_{\Lambda_i}$. 

{\it Remark 2}. The asymptotic distributions (\ref{d*1}) and (\ref{d*2}) admit  heavy tails.
In the case (ii) we obtain a power law asymptotic degree distribution (\ref{d*1}) provided that at least one 
of $P_1$ and $P_2$ has a power law and $P_1(0),P_2(0)<1$.
In the case (iii) 
we obtain a power law asymptotic degree distribution (\ref{d*2}) provided that  $P_2$ has a power law.

{\it Remark 3}. Since the second moment $a_2$ does not show up in (\ref{d*1}), (\ref{d*1++}) we expect that in the case 
(ii) the second moment condition $\E X_1^2<\infty$ is redundant and  could perhaps be replaced by the weaker first moment condition $\E X_1<\infty$.

 Random intersection graphs 
 have 
 attracted considerable attention in the recent literature, see, e.g., 
 \cite{Barbour2011},
 \cite{behrisch2007},
 \cite{Blackburn2009},
 \cite{Bradonjic2010},
 \cite{Britton2008},
 \cite{eschenauer2002},
 \cite{Rybarczyk2011}. 
 Starting with the paper  
  by Karo\'nski et al \cite{karonski1999}, see also \cite{Singer1995},
  where the  case  of degenerate distributions $P_1=P_{1n}$, $P_2=P_{2n}$ depending on $n$ was considered 
  (i.e., $\PP_{1n}(c_n)=\PP_{2n}(c_n)=1$,  for some $c_n>0$), 
 several more complex random intersection graph models were later introduced by Godehardt and Jaworski \cite{godehardt2001}, Spirakis et al. \cite{Spirakis2004}, Shang \cite{Shang2010}.  
%Guillaume and  Latapy
%\cite{Guillaume+L2004} argue that many complex networks expoit an underlying
%bipartite graph structure, like the bipartite graph $H$ above.
%Independently, Eshenauer and Gligor \cite{eschenauer2002} came up with a  random intersection graph  as a model of secure wireless %sensor network that uses random predistribution of keys. 
 The asymptotic degree distribution  for various random intersection graph models was shown in  
  \cite{Bloznelis2008}, \cite{Bloznelis2010}, \cite{Bloznelis2010+}, \cite{Bloznelis2011+},  \cite{Deijfen}, \cite{JKS}, \cite{JaworskiStark2008},
  \cite{Rybarczyk-degree2011}, \cite{stark2004}.

\section{Proofs}

Before the proof we introduce some notation and give two auxiliary lemmas.

The event that the edge $\{w_i, v_j\}$ is present in $H=H_{X,Y}$ is denoted $w_i\to v_j$. We denote
\begin{displaymath}
 {\mathbb I}_{ij}={\mathbb I}_{\{w_i\to v_j\}},
\qquad
{\mathbb I}_i={\mathbb I}_{i1},
\qquad
u_i=\sum_{2\le j\le n}{\mathbb I}_{ij},
\qquad
L=\sum_{i=1}^mu_i{\mathbb I}_i.
\end{displaymath}
We remark, that $u_i$ counts all neighbours of $w_i$ in $H$ 
belonging to the set $V\setminus \{v_1\}$. Denote
\begin{eqnarray}\nonumber
 &&
{\hat a}_k=m^{-1}\sum_{1\le i\le m}X^k_i,
\qquad
{\hat b}_k=n^{-1}\sum_{2\le j\le n}Y^k_j,
\\
\label{S-XY}
&&
\lambda_{ij}=\frac{X_iY_j}{\sqrt{mn}},
\qquad
 S_{XY}=\sum_{i=1}^m\sum_{j=2}^n\frac{X_i}{\sqrt{nm}}\lambda_{ij}\min\{1,\lambda_{ij}\},
\end{eqnarray}
and introduce the event ${\cal A}_1=\{\lambda_{i1}<1, \, 1\le i\le m\}$.
By ${\PP}_1$ and $\E_1$ we denote the conditional probability and conditional expectation given $Y_1$.
By ${\tilde \PP}$ and ${\tilde \E}$ we denote the conditional probability and conditional expectation 
given $X,Y$.
By  $d_{TV}(\zeta,\xi)$ we denote the total variantion 
distance between the probability distributions
of
 random variables $\zeta$ and $\xi$. In the case where $\zeta,\xi$ and $X, Y$ are defined on the same probability space we denote 
 by 
 ${\tilde d}_{TV}(\zeta,\xi)$ the total 
variation distance between the conditional distributions of $\zeta$ and $\xi$ given $X,Y$.

In the proof  below we use the following 
simple fact about the convergence of a sequence of random variables $\{\varkappa_n\}$: 
\begin{equation}\label{fact}
 \varkappa_n=o_P(1),
\quad
\E \sup_n\varkappa_n<\infty
\quad
\Rightarrow
\quad
\E\varkappa_n=o(1).
\end{equation}
We remark that the condition $\E \sup_n\varkappa_n<\infty$ (which assumes implicitly that
all $\varkappa_n$ are defined on the same probability space) can be replaced by the more 
resctrictive condition that there exists a constant $c>0$ such that 
$|\varkappa_n|\le c$, for all $n$. In the latter case we do not need all $\varkappa_n$ to be defined on the same probability space.
In particular, given a sequence of bivariate random vectors $\{(\eta_n,\theta_n)\}$ such that, for  every $n$ and $m=m_n$ random variables $\eta_n$, $\theta_n$ and
 $\{X_i\}_{i=1}^m$, $\{Y_j\}_{j=1}^n$ are defined on the same probability space, we have 
\begin{displaymath}
{\tilde d}_{TV}(\eta_n,\theta_n)=o_P(1)
\
\Rightarrow
\
d_{TV}(\eta_n,\theta_n)\le \E{\tilde d}_{TV}(\eta_n,\theta_n) =o(1).
\end{displaymath}

\begin{lem}\label{d-L} Assume that $\E X_1^2<\infty$ and $\E Y_1<\infty$.
We have as $n,m\to+\infty$
\begin{eqnarray}\label{d-L21}
&&
\PP(d(v_1)\not=L)=o(1),
\\
\label{d-L22}
&&
\PP({\cal A}_1)=1-o(1),
\\
\label{d-L23}
&&
S_{XY}=o_P(1),
\qquad
\E S_{XY}=o(1).
\end{eqnarray}

\end{lem}
\begin{proof}[Proof of Lemma \ref{d-L}] Proof of (\ref{d-L21}).
We observe that the event ${\cal A}=\{d(v_1)\not=L\}$ occurs in the case where for some $2\le j\le n$
and some distinct $i_1,i_2\in [m]$ the event 
${\cal A}_{i_1,i_2,j}=
\{w_{i_1}\to v_1,\, w_{i_1}\to v_j,\, w_{i_2}\to v_1,\, w_{i_2}\to v_j\}$ occurs.
From the union bound and the inequality 
${\tilde \PP}({\cal A}_{i_1,i_2,j})\le m^{-2}n^{-2}X_{i_1}^2X_{i_2}^2Y_1^2Y_j^2$ we obtain
\begin{equation}\label{d-L1}
{\tilde \PP}({\cal A})
=
{\tilde \PP}
\left(\bigcup_{\{i_1, i_2\}\subset[m]}\bigcup_{2\le j\le n}{\cal A}_{i_1,i_2,j}\right)
\le n^{-1}{\hat b}_2Y_1^2 Q_X.
\end{equation}
Here $Q_X=m^{-2}\sum_{\{i_1,i_2\}\subset[m]}X_{i_1}^2X_{i_2}^2$. We note that
$Q_X$ and  $Y_1^2$ are stochastically bounded  and 
%We note that the product $Y_1^2Q_X$ is stochastically bounded and 
$n^{-1}{\hat b}_2=o_P(1)$ 
%converges to zero in probability
as $n\to+\infty$. Therefore, 
%the random variable 
${\tilde \PP}({\cal A})=o_P(1)$. 
%converges to zero in probability. 
Now (\ref{fact}) implies (\ref{d-L21}).
%This implies (\ref{d-L21}) since, for any $\varepsilon\in (0,1)$,
%\begin{displaymath}
%\PP({\cal A})=\E ({\tilde \PP}({\cal A}))
%\le 
%\varepsilon+\PP({\tilde \PP}({\cal A})>\varepsilon)
%=\varepsilon+o(1).
%\end{displaymath}

\medskip

Proof of (\ref{d-L22}).
Let ${\overline {\cal A}}_1$ denote the complement event to ${\cal A}_1$.
%Next we show that $\PP({\overline{\cal A}}_1)=o(1)$. 
%We obtain from 
By 
the union bound  and Markov's inequality 
%that
\begin{displaymath}
\PP_1({\overline {\cal A}}_1)\le \sum_{i\in [m]}\PP_1(\lambda_{i1}\ge 1)
\le 
(nm)^{-1}Y_1^2\sum_{i\in [m]}\E X_i^2=n^{-1}a_2Y_1^2.
\end{displaymath}
Hence we obtain $\PP_1({\overline {\cal A}}_1)=o_P(1)$.  Now (\ref{fact}) implies $\PP({\overline {\cal A}}_1)=\E \PP_1({\overline {\cal A}}_1)=o(1)$.
%(\ref{d-L22}).
%Note that this implies (\ref{d-L22}) since, 
%for any $\varepsilon\in (0,1)$, we have
%\begin{displaymath}
% \PP({\overline{\cal A}}_1)
%=
%\E\PP_1({\overline {\cal A}}_1)
%\le 
%\varepsilon+\PP(\PP_1({\overline {\cal A}}_1)>\varepsilon)
%=
%\varepsilon+o(1).
%\end{displaymath}

\medskip

Proof of (\ref{d-L23}). Since the first bound of (\ref{d-L23}) follows from the second one, we only prove the latter.
%Now, we show that $S_{XY}=o_P(1)$.
Denote ${\hat X}_1=\max\{X_1, 1\}$ and ${\hat Y}_1=\max\{Y_1,1\}$.
We observe that $\E X_1^2<\infty$, $\E Y_1<\infty$ implies
\begin{displaymath}
\lim_{t\to+\infty}\E {\hat X}_1^2{\mathbb I}_{\{{\hat X}_1>t\}}=0,
\qquad
\lim_{t\to+\infty}\E {\hat Y}_1{\mathbb I}_{\{{\hat Y}_1>t\}}=0. 
\end{displaymath}
 Hence one can find
a strictly increasing function $\varphi:[1,+\infty)\to [0,+\infty)$ 
with $\lim_{t\to+\infty}\varphi(t)=+\infty$
such that
\begin{equation}\label{hat1}
\E {\hat X}_1^2\varphi({\hat X}_1)<\infty,
\qquad  
\E {\hat Y}_1\varphi({\hat Y}_1)<\infty.
\end{equation} 
In addition, we can choose $\varphi$  satisfying 
\begin{equation}\label{phi}
\varphi(t)<t
\qquad
{\text{and}}
\qquad
\varphi(st)
\le \varphi(t)\varphi(s), 
\qquad
\forall s,t\ge 1. 
\end{equation}
For this purpose we take a sufficiently slowly growing concave function 
$\psi:[0,+\infty)\to [0,+\infty)$ with $\psi(0)=0$ and define
$\varphi(t)=e^{\psi(\ln (t))}$. We note that the second inequality of (\ref{phi})
follows from the concavity property of $\psi$. We remark, that (\ref{phi})
implies
\begin{equation}\label{phi1}
 t/(st)=s^{-1}\le 1/\varphi(s)\le\varphi(t)/\varphi(st),
\qquad
s,t\ge 1.
\end{equation}

Let ${\hat S}_{XY}$ be defined as in  (\ref{S-XY}) above, but with
$\lambda_{ij}$ replaced by ${\hat\lambda}_{ij}={\hat X}_i{\hat Y}_j/\sqrt{mn}$.
We note, that $S_{XY}\le {\hat S}_{XY}$. Furthermore, from the inequalities 
\begin{equation}\label{phi2}
 \min\{1,{\hat\lambda}_{ij}\}
\le
\min\Bigl\{1,\frac{\varphi({\hat X}_i{\hat Y}_j)}{\varphi(\sqrt{mn})}\Bigr\}
\le
\frac{\varphi({\hat X}_i{\hat Y}_j)}{\varphi(\sqrt{mn})}
\le 
\frac{\varphi({\hat X}_i)\varphi({\hat Y}_j)}{\varphi(\sqrt{mn})}
\end{equation}
we obtain 
$S_{XY}\le {\hat S}_{XY}\le S^*_{XY}/\varphi(\sqrt{nm})$, 
where 
\begin{displaymath}
S^*_{XY}
=
(mn)^{-1}
\left(\sum_{1\le i\le m}{\hat X}^2_i\varphi({\hat X}_i)\right)
\left(\sum_{2\le j\le n}{\hat Y}_j\varphi({\hat Y}_j)\right).
\end{displaymath}
We remark that (\ref{phi}) and (\ref{phi1}) and  imply the third and the first inequality of 
(\ref{phi2}), respectively. 

Finally, the bound $\E S_{XY}=o(1)$ follows from the inequality 
$S_{XY}\le S^*_{XY}/\varphi(\sqrt{nm})$
and the fact  that $\E S^*_{XY}$ remains bounded as $n,m\to+\infty$, 
see  (\ref{hat1}).
\end{proof}

%*****************************************************************************************************************************

In the proof of Theorem \ref{T1} we use the following inequality 
 refered to as LeCam's lemma, see e.g., \cite{Steele}.
\begin{lem}\label{LeCamLemma} Let $S={\mathbb I}_1+ {\mathbb I}_2+\dots+ {\mathbb I}_n$ be the sum of independent random indicators
with probabilities $\PP({\mathbb I}_i=1)=p_i$. Let $\Lambda$ be Poisson random variable with mean $p_1+\dots+p_n$. The total variation
distance between the distributions $P_S$ of $P_{\Lambda}$ of $S$ and $\Lambda$
\begin{equation}\label{LeCam}
\sup_{A\subset \{0,1,2\dots \}}|\PP(S\in A)-\PP(\Lambda\in A)|= \frac{1}{2} \sum_{k\ge 0}|\PP(S=k)-\PP(\Lambda=k)|
\le
\sum_{i}p_i^2.
\end{equation}
\end{lem}

%*****************************************************************************************************************************

\begin{proof}[Proof of Theorem \ref{T1}]

{\it The case (i).} We have $\PP(d(v_1)>\varepsilon)\le \PP(L>\varepsilon)$, for each $\varepsilon>0$. We prove that $\PP(L>\varepsilon)=o(1)$, for any $\varepsilon>0$.
In view of the identity $\PP(L>\varepsilon)=\E\PP_1(L>\varepsilon)$ and (\ref{fact}) it suffices to show that $\PP_1(L>\varepsilon)=o_P(1)$.
For this purpose we write, by the union bound and Markov's inequality,  
\begin{displaymath}
\PP_1(L>\varepsilon)
\le 
\sum_{1\le k\le m}\PP_1({\mathbb I}_k=1)
\le
\E_1\sum_{1\le k\le m}\lambda_{k1}
=
\sqrt{m/n}Y_1\E X_1
=
o_P(1).
\end{displaymath}

In cases (ii) and (iii) we apply (\ref{d-L21}). In view of 
(\ref{d-L21}) the random variables $d(v_1)$ and $L$ have the same asymptotic distribution (if any).
Hence, it suffices to
%We will 
show 
the convergence in distribution of $L$.

{\it The case (ii)}.
Here we prove that  $L$ converges in distribution to (\ref{d*1}). 
 We first approximate $L$ by 
the  random variable $L_3=\sum_{k=1}^m\eta_k\xi_{3k}$. 
Then we show that $L_3$ converges in distribution to (\ref{d*1}).
Here $\eta_1,\dots, \eta_m$, $\xi_{3 1},\dots, \xi_{3 m}$ are 
conditionally independent (given $X,Y$) 
Poisson random variables with ${\tilde \E}\eta_k=\lambda_{k1}$ and 
${\tilde \E}\xi_{3 k}=X_k(n/m)^{1/2}b_1$. We assume, in addition, 
that given $X,Y$, the 
sequences $\{{\mathbb I}_{k}\}_{k=1}^m$ and $\{\xi_{3k}\}_{k=1}^m$ 
are conditionally independent.
 
Given $X,Y$, we generate independent Poisson random variables
$\xi_{11},\dots, \xi_{1m}$, $\Delta_{11},\dots, \Delta_{1m}$, with the conditional mean values
\begin{displaymath}
{\tilde \E}\xi_{1k}= \sum_{2\le j\le n}p_{kj},
\qquad 
{\tilde \E}\Delta_{1k}=\sum_{2\le j\le n}(\lambda_{kj}-p_{kj}),
\qquad
1\le k\le m.
\end{displaymath}
We assume that, given $X,Y$, these Poisson random variables are conditionally independent of the sequence 
$\{\eta_k\}_{k=1}^m$. We suppose, in addition, that $\{\eta_k\}_{k=1}^m$ is
conditionally  independent (given $X, Y_1$)  of the set of edges of $H$ 
that are not incident to 
$v_1$. 
We define  $\xi_{2k}=\xi_{1k}+\Delta_{1k}$ and observe that $\xi_{2k}$ has conditional (given $X,Y$) Poisson distribution with
the conditional mean value ${\tilde \E}{\xi}_{2k}=\sum_{2\le j\le n}\lambda_{kj}$.
Introduce the random variables 
\begin{displaymath}
L_0=\sum_{1\le k\le m}\eta_ku_k,
\qquad
L_1=\sum_{1\le k\le m}\eta_k\xi_{1k},
\qquad
L_2=\sum_{1\le k\le m}\eta_k\xi_{2k}.
\end{displaymath}
In order to show that $L$ and $L_3$ 
have the same asymptotic probability distribution (if any) we 
prove that 
\begin{eqnarray}\label{DeltaT}
&&
d_{TV}(L, L_0)=o(1),
\qquad
d_{TV}(L_0,L_1)=o(1),
\\
\label{L-L}
&&
\E|L_1-L_2|=o(1),
\qquad
\qquad
{\tilde L}_2-{\tilde L}_3=o_P(1).
\end{eqnarray}
Here ${\tilde L}_2$ and ${\tilde L}_3$ are marginals of the random 
vector $({\tilde L}_2, {\tilde L}_3)$ constructed below which has the 
property that ${\tilde L}_2$ has the same distribution as $L_2$ and ${\tilde L}_3$ has 
the same distribution as $L_3$.

Let us prove the first bound of (\ref{DeltaT}).
In view of (\ref{fact}) it suffices to show that
${\tilde d}_{TV}(L_0,L)=o_P(1)$. In order to prove the latter bound we apply the inequality
\begin{equation}\label{TV1}
 {\tilde d}_{TV}(L_0, L){\mathbb I}_{{\cal A}_1}\le n^{-1}Y_1^2{\hat a}_2
\end{equation} 
shown below. 
We remark that (\ref{TV1}) implies 
\begin{displaymath}
 {\tilde d}_{TV}(L_0,L)
 \le 
 {\tilde d}_{TV}(L_0,L) {\mathbb I}_{{\cal A}_1}+{\mathbb I}_{{\overline{\cal A}}_1}
 \le
 n^{-1}Y_1^2{\hat a}_2+{\mathbb I}_{{\overline{\cal A}}_1}=o_P(1).
\end{displaymath}
Here $n^{-1}Y_1^2{\hat a}_2=o_P(1)$, because $Y_1^2{\hat a}_2$ is stochastically bounded. Furthermore,  the bound
${\mathbb I}_{{\overline{\cal A}}_1}=o_P(1)$
follows from (\ref{d-L22}).

It remains to prove (\ref{TV1}).
% from LeCam's inequality, see Lemma \ref{LeCamLemma} below. 
We denote  
$L'_k=\sum_{i=1}^k{\mathbb I}_iu_i+\sum_{i=k+1}^m\eta_iu_i$ and write, by the triangle inequality, 
%of the total variation distance,
\begin{displaymath}
 {\tilde d}_{TV}(L_0,L)\le \sum_{k=1}^m{\tilde d}_{TV}(L'_{k-1}, L'_k).
\end{displaymath}
Then we estimate 
${\tilde d}_{TV}(L'_{k-1}, L'_k)
\le 
{\tilde d}_{TV}(\eta_k,{\mathbb I}_k)\le (nm)^{-1}Y_1^2X_k^2$.
Here the first inequality follows from the properties of the total variation distance. 
The second inequality follows from Lemma \ref{LeCamLemma} and the fact 
that on the event ${\cal A}_1$ we have
$p_{k1}=\lambda_{k1}$.

\medskip

Let us prove the second bound of (\ref{DeltaT}).
In view of (\ref{fact}) it suffices to  show that
  ${\tilde d}_{TV}(L_0,L_1)=o_P(1)$. We denote  
$L^*_k=\sum_{i=1}^k\eta_iu_i+\sum_{i=k+1}^m\eta_i\xi_{1i}$ and write, 
by the triangle inequality, 
%of the total variation distance,
\begin{equation}\label{dTV4}
 {\tilde d}_{TV}(L_0,L_1)\le \sum_{k=1}^m{\tilde d}_{TV}(L^*_{k-1}, L^*_k).
\end{equation}
Here
\begin{equation}\label{dTV41}
{\tilde d}_{TV}(L^*_{k-1}, L^*_k)\le {\tilde d}_{TV}(\eta_ku_k,\eta_k\xi_{1k})
\le 
{\tilde \PP}(\eta_k\not=0){\tilde d}_{TV}(u_k,\xi_{1k}).
\end{equation}
Now, invoking the inequalities  
${\tilde \PP}(\eta_k\not=0)=1-e^{-\lambda_{k1}}\le \lambda_{k1}$
and ${\tilde d}_{TV}(u_k,\xi_{1k})\le \sum_{j=2}^np^2_{kj}$, we obtain from 
(\ref{dTV4}), (\ref{dTV41})  and (\ref{d-L23}) that
\begin{displaymath}
 {\tilde d}_{TV}(L_0,L_1)\le \sum_{k=1}^m\lambda_{k1}\sum_{j=2}^np^2_{kj} 
\le 
Y_1S_{XY}=o_P(1).
\end{displaymath}

\medskip

Let us prove the first  bound of (\ref{L-L}).
We observe that 
\begin{displaymath}
 |L_2-L_1|=L_2-L_1
=
\sum_{1\le k\le m}\eta_k\Delta_{1k}
\end{displaymath}
and 
\begin{displaymath}
{\tilde \E}\sum_{1\le k\le m}\eta_k\Delta_{1k}
=
\sum_{1\le k\le m}\lambda_{k1}\sum_{2\le j\le n}
(\lambda_{kj}-1){\mathbb I}_{\{\lambda_{kj}>1\}}
\le Y_1S_{XY}. 
\end{displaymath}
We obtain $\E|L_2-L_1|\le \E Y_1\E S_{XY}=o(1)$, see
 (\ref{d-L23}).

\medskip

Let us prove the second  bound of (\ref{L-L}).  Given $X,Y$, generate 
independent Poisson random variables
$\xi'_{31},\dots,\xi'_{3m}, \Delta_{21},\dots,\Delta_{2m}, \Delta_{31},\dots, \Delta_{3m}$ which are conditionally independent of the sequence $\{\eta_k\}_{k=1}^m$ and have  the conditional mean values
\begin{displaymath}
{\tilde \E}\xi'_{3k}=X_k(n/m)^{1/2}b,
\qquad
{\tilde \E}\Delta_{2k}=X_k(n/m)^{1/2}\delta_{2},
\qquad
{\tilde \E}\Delta_{3k}=X_k(n/m)^{1/2}\delta_{3}.
\end{displaymath}
Here
$b=\min\{{\hat b}_1, b_1\}$, $\delta_{2}= {\hat b}_1-b$, $\delta_{3}=b_1-b$. We note that $\delta_2,\delta_3\ge 0$ and
observe that the random vector
\begin{displaymath}
(L'_2,L'_3),
\qquad
L'_2=\sum_{1\le k\le m}\eta_k(\xi'_{3k}+\Delta_{2k}), 
\qquad
L'_3=\sum_{1\le k\le m}\eta_k(\xi'_{3k}+\Delta_{3k})
\end{displaymath}
has the marginal distributions of $(L_2,L_3)$. In addition, we have 
\begin{equation}\label{L2-L3}
 {\tilde \E}|L'_2-L'_3|
=
(\delta_2+\delta_3)Y_1{\hat a}_2=|{\hat b}_1-b_1|Y_1{\hat a}_2=o_P(1).
\end{equation}
In the last step we used the fact that $Y_1{\hat a}_2=O_P(1)$ and ${\hat b}_1-b_1=o_P(1)$, 
by the law of large numbers.  
Finally, we show that (\ref{L2-L3}) implies
the bound $|L'_2-L'_3|=o_P(1)$. Denoting, for short, $H=|L_2'-L_3'|$ and 
$h={\tilde \E}{\mathbb I}_{\{H\ge \varepsilon\}}$ we write, for $\varepsilon\in (0,1)$,
\begin{equation}\label{HHHH}
 \PP(H\ge\varepsilon)
=
\E h
=
\E h
({\mathbb I}_{\{{\tilde \E}H\ge\varepsilon^2\}}
+
{\mathbb I}_{\{{\tilde \E}H<\varepsilon^2\}}).
\end{equation}
Using the simple inequality $h\le 1$ and the inequality,
$h\le \varepsilon^{-1}{\tilde \E}H$, which follows from Markov's inequality, we obtain
\begin{eqnarray}\nonumber
 \E h
{\mathbb I}_{\{{\tilde \E}H\ge\varepsilon^2\}}
&
\le 
&
\E 
{\mathbb I}_{\{{\tilde \E}H\ge\varepsilon^2\}}=\PP({\tilde \E}H\ge\varepsilon^2)=o(1),
\\
\nonumber
\E h {\mathbb I}_{\{{\tilde \E}H<\varepsilon^2\}}
&
\le 
&
\E (\varepsilon^{-1}{\tilde \E}H){\mathbb I}_{\{{\tilde \E}H<\varepsilon^2\}}
< 
\varepsilon.
\end{eqnarray}
Invoking these inequalities in (\ref{HHHH}) we obtain $\PP(H\ge \varepsilon)<\varepsilon+o(1)$.
Hence $H=o_P(1)$. 

% Indeed, denote
% $H:=|{\tilde L}'-{\bar L}'|$, ${\tilde H}:={\tilde E}H$ and write for 
%any $\varepsilon\in (0,1)$ 
%\begin{displaymath}
% \PP(H>\varepsilon)\le\PP({\tilde H}>\varepsilon^2)
%+
%\PP(H>\varepsilon, {\tilde H}\le\varepsilon^2).
%\end{displaymath}
%Next we apply Markov's inequality to the second probability
%\begin{displaymath}
% \PP(H>\varepsilon, {\tilde H}\le\varepsilon^2)
%=
%\E{\mathbb I}_{\{ H>\varepsilon \}} {\mathbb I}_{\{ {\tilde H}\le \varepsilon^2\}}
%\le
%\varepsilon^{-1}\E H{\mathbb I}_{\{{\tilde H}\le \varepsilon^2\}}
%\end{displaymath}
%and then invoke the simple inequality
%\begin{displaymath}
% \E (H{\mathbb I}_{\{{\tilde H}\le \varepsilon^2\}})
%=\E({\tilde \E} (H{\mathbb I}_{\{{\tilde H}\le \varepsilon^2\}}))=
%\E ({\mathbb I}_{\{{\tilde H}\le \varepsilon^2\}} {\tilde H})
%\le \varepsilon^2.
%\end{displaymath}
%We obtain $\PP(H>\varepsilon)\le\PP({\tilde H}>\varepsilon^2)+\varepsilon=o(1)+\varepsilon$,
%thus proving the bound $H=o_P(1)$.

Now we prove that  $L_3$ converges in distribution to (\ref{d*1}).
 Introduce the random variable ${\bar L}=\sum_{1\le k\le m}\eta_k{\bar \xi}_k$, where,
given $X,Y$, the random variables ${\bar \xi}_1,\dots, {\bar \xi}_m$ 
are conditionally independent of $\{\eta_k\}_{k=1}^m$ and have the 
conditional mean values ${\tilde \E}{\bar \xi}_k=X_k\beta^{-1/2}b_1$. 
Proceeding as in the proof of the second bound of 
(\ref{L-L}) above, we construct a random vector 
$(L_3'', {\bar L}')$ 
with the same marginals as $(L_3, {\bar L})$ 
and such that 
\begin{equation}\label{LLL}
{\tilde \E}|L''_3- {\bar L}'|
\le 
|1-(m/n)^{1/2}\beta^{-1/2}|Y_1b_1{\hat a}_2=o_P(1).
\end{equation}
In the last step we used the fact that $Y_1{\hat a}_2=O_P(1)$ and $m/n\to\beta$. 
Now, (\ref{LLL}) implies $L''_3-{\bar L}'=o_P(1)$. 
We conclude that $L_3$ and ${\bar L}$ have the same asymptotic distribution
(if any). 

Next  we prove that ${\bar L}$ converges in distribution to (\ref{d*1}). For this purpose we 
 show that  $\E e^{it{\bar L}}\to\E e^{itd_*}$, 
for each $t\in (-\infty,+\infty)$. Denote 
$\Delta(t)=e^{it{\bar L}}-e^{itd_*}$. We shall show below 
that, for any real $t$ and any realized value $Y_1$ 
there exists a positive constant 
$c=c(t,Y_1)$ such that for every $0<\varepsilon<0.5$ 
we have
\begin{equation}\label{epsilon}
 \limsup_{n,m\to+\infty}|\E(\Delta(t)|Y_1)|<c\varepsilon.
\end{equation}
Clearly, (\ref{epsilon}) implies $\E(\Delta(t)| Y_1)=o(1)$.
 This fact together with the simple inequality
 $|\Delta(t)|\le 2$  yields $\E\Delta(t)=o(1)$, 
by Lebesgue's dominated convergence theorem.
Observing that  $\E\Delta(t)=\E e^{it{\bar L}}-\E e^{itd_*}$ we conclude that
$\E e^{it{\bar L}}\to\E e^{itd_*}$.

\bigskip

We fix $0<\varepsilon<0.5$ and prove (\ref{epsilon}). 
Before the proof we introduce  some notation. Denote
\begin{eqnarray}\nonumber
 &&
f_\tau(t)=\E e^{it\tau_1},
\qquad
{\bar f}_{\tau}(t)=\sum_{r\ge 0}e^{itr}{\bar p}_r, 
\qquad
{\bar p}_r
=
{\bar \lambda}^{-1}\sum_{1\le k\le m}\lambda_{k1}{\mathbb I}_{\{{\bar\xi}_k=r\}},
\qquad
{\bar \lambda}=\sum_{k=1}^m\lambda_{k1},
\\
\nonumber
&&
\delta=({\bar f}_\tau(t)-1){\bar \lambda}-(f_{\tau}(t)-1)\lambda_1,
\qquad
f(t)=\E_1e^{itd_*}, 
\qquad
{\bar f}(t)={\bar\E}e^{it{\bar L}}.
\end{eqnarray}
Here ${\bar \E}$ denotes the conditional expectation given $X,Y$ 
and ${\bar\xi}_1,\dots, {\bar\xi}_m$.
Introduce the event
${\cal D}=\{|{\hat a}_1-a_1|<\varepsilon\min\{1, a_1\}\}$ and let ${\overline{\cal D}}$ denote the complement event.  
Furthermore, select the number 
$T>1/\varepsilon$  such that $\PP(\tau_1\ge T)< \varepsilon$. By $c_1, c_2, \dots$ 
we denote positive numbers which do not depend on $n,m$.

We observe that, given $Y_1$, the conditional distribution of $d_*$ is the 
compound Poisson distribution
with the characteristic function $f(t)=e^{\lambda_1(f_{\tau}(t)-1)}$. Similarly,
given $X,Y$ 
and ${\bar\xi}_1,\dots, {\bar\xi}_m$, 
the conditional distribution of ${\bar L}$ is the 
compound Poisson distribution
with the characteristic function ${\bar f}(t)=e^{{\bar \lambda}({\bar f}_{\tau}(t)-1)}$.
In the proof of (\ref{epsilon}) we exploit the convergence
${\bar\lambda}\to\lambda_1$ and ${\bar f}_{\tau}(t)\to f_{\tau}(t)$. In what follows 
we assume that $m,n$ are so 
large that $\beta\le 2m/n\le 4\beta$.

Let us prove  (\ref{epsilon}). 
We write 
\begin{displaymath}
\E_1\Delta(t)=I_{1}+I_{2},
\qquad
I_{1}=\E_1\Delta(t){\mathbb I}_{\cal D},
\qquad
I_2=\E_1\Delta(t){\mathbb I}_{\overline{\cal D}}.         
\end{displaymath}
Here $|I_2|\le 2\PP_1({\overline {\cal D}})=2\PP({\overline {\cal D}})=o(1)$, by the law of large numbers.
Next we estimate $I_{1}$. Combining the identity $\E_1\Delta(t)=\E_1f(t)(e^{\delta}-1)$
with the inequalities $|f(t)|\le 1$ and $|e^{s}-1|\le |s|e^{|s|}$, 
we obtain
\begin{equation}
 |I_1|\le \E_1|\delta|e^{|\delta|}{\mathbb I}_{\cal D}
\le 
c_1\E_1|\delta|{\mathbb I}_{\cal D}.
\end{equation}
Here we estimated $e^{|\delta|}\le e^{8\lambda_1}=:c_1$ using the inequalities
 \begin{displaymath}
 |\delta|\le 2{\bar \lambda}+2\lambda_1,
\qquad
{\bar \lambda}=Y_1(m/n)^{1/2}{\hat a}_1\le 3\lambda_1. 
 \end{displaymath}
We remark that the last  inequality holds for 
$m/n\le 2\beta$ provided that the event ${\cal D}$ occurs.
Finally, we show that $\E_1|\delta|{\mathbb I}_{\cal D}\le (c_2+\lambda_1c_3+\lambda_1c_4)\varepsilon+o(1)$. 
We first write
\begin{displaymath}
\delta
=
({\bar f}_{\tau}(t)-1)({\bar\lambda}-\lambda_1)+({\bar f}_{\tau}(t)-f_{\tau}(t))\lambda_1,
\end{displaymath}
and estimate $|\delta|\le 2|{\bar\lambda}-\lambda_1|+\lambda_1|{\bar f}_{\tau}(t)-f_{\tau}(t)|$.
From the inequalities $|{\hat a}_1-a_1|<\varepsilon$ and $m/n\le 2\beta$ we obtain
\begin{displaymath}
 |{\bar\lambda}-\lambda_1|
\le 
Y_1|{\hat a}_1-a|(m/n)^{1/2}
+Y_1a_1|(m/n)^{1/2}-\beta^{1/2}|
\le 
2Y_1\beta^{1/2}\varepsilon+o(1).
\end{displaymath}
Hence $\E_1|{\bar\lambda}-\lambda_1|{\mathbb I}_{\cal D}\le c_2\varepsilon+o(1)$,
where $c_2=2Y_1\beta^{1/2}$.
We secondly  show that
\begin{displaymath}
 \E_1|{\bar f}_{\tau}(t)-f_{\tau}(t)|{\mathbb I}_{\cal D}\le (c_3+c_4)\varepsilon+o(1).
\end{displaymath}
To this aim we  split 
\begin{displaymath}
 {\bar f}_{\tau}(t)-f_{\tau}(t)=\sum_{r\ge 0}e^{itr}({\bar p}_r-p_r)=R_1-R_2+R_3,
\end{displaymath}
and estimate separately the terms
\begin{displaymath}
 R_1=\sum_{r\ge T}e^{itr}{\bar p}_r,
\qquad
R_2=\sum_{r\ge T}e^{itr}p_r,
\qquad 
R_3=\sum_{0\le r<T}e^{itr}({\bar p}_r-p_r).
\end{displaymath}
Here we denote
$p_r=\PP(\tau_1=r)$. The upper bound for $R_2$ follows by the choice of $T$ 
\begin{displaymath}
 |R_2|\le \sum_{r\ge T}p_r=\PP(\tau_1\ge T)< \varepsilon.
\end{displaymath}
Next,
combining the identity
${\bar p}_r=({\hat a}_1m)^{-1}\sum_{1\le k\le m}X_k{\mathbb I}_{\{{\bar \xi}_k=r\}}$
with the inequalities
\begin{eqnarray}
 |R_1|
\le 
({\hat a}_1m)^{-1}
\sum_{r\ge T}\sum_{1\le k\le m}X_k{\mathbb I}_{\{{\bar \xi}_k=r\}}
=
({\hat a}_1m)^{-1}\sum_{1\le k\le m}X_k{\mathbb I}_{\{{\bar \xi}_k\ge T\}}
\end{eqnarray}
and ${\hat a}_1^{-1}{\mathbb I}_{\cal D}\le 2a_1^{-1}$, we estimate
\begin{equation}\nonumber
 \E_1|R_1|{\mathbb I}_{\cal D}\le 2a_1^{-1}\E_1(X_1{\mathbb I}_{\{{\bar \xi}_1\ge T\}})
\le 
2a_1^{-1}T^{-1}\E_1(X_1{\bar\xi}_1)
=
2a_1^{-1}a_2b_1\beta^{-1/2}\varepsilon.
\end{equation}
Hence $ \E_1|R_1|{\mathbb I}_{\cal D}\le c_4\varepsilon$, where $c_4=2a_1^{-1}a_2b_1\beta^{-1/2}$.

Now we estimate $R_3$. 
We denote $p'_r=({\hat a}_1/a_1){\bar p}_r$
and observe that the inequality $|{\hat a}_1-a_1|\le \varepsilon a_1$ implies
$|{\hat a}_1a_1^{-1}-1|\le \varepsilon$ and
\begin{displaymath}
 |\sum_{0\le r\le T}e^{itr}({\bar p}_r- p'_r)|\le \varepsilon\sum_{0\le r\le T}
{\bar p}_r\le \varepsilon.
\end{displaymath}
In the last inequality we use the fact that the probabilities  $\{{\bar p}_r\}_{r\ge 0}$ 
sum up to $1$. It follows now that
\begin{displaymath}
 |R_3|{\mathbb I}_{\cal D}\le \varepsilon+\sum_{0\le r\le T}|p'_r-p_r|.
\end{displaymath}
Furthermore, observing that 
$\E_1p'_r=a^{-1}\E_1X_k{\mathbb I}_{\{{\bar\xi}_k=r\}}=p_r$, for $1\le k\le m$,
 we obtain 
\begin{displaymath}
 \E_1|p'_r-p_r|^2
= 
m^{-1}\E_1|a^{-1}X_1{\mathbb I}_{\{{\bar\xi}_1=r\}}-p_r|^2\le m^{-1}a_1^{-2}\E X_1^2.
\end{displaymath}
Hence, $\E_1|p'_r-p_r|=O(m^{-1/2})$. We conclude that
\begin{displaymath}
\E_1 |R_3|{\mathbb I}_{\cal D}
\le 
\varepsilon+O(|T|m^{-1/2})=\varepsilon+o(1).
\end{displaymath}

{\it The case (iii)}.  We start with introducing some notation.
Denote $m/n=\beta_n$. Given $\varepsilon\in (0,1)$ 
introduce random variables
% ${\mathbb I}'_{k\varepsilon}={\mathbb I}_{\{X_k(n/m)^{1/2}b_1<\varepsilon\}}.
\begin{displaymath}
\gamma=\sum_{1\le k\le m}{\mathbb I}'_{k}\lambda_{k1}\gamma_{k},
\qquad
\gamma_{k}=X_k\beta_n^{-1/2}b_1{\mathbb I}'_{k},
\qquad
{\mathbb I}'_{k}={\mathbb I}_{\{X_k\beta_n^{-1/2}b_1<\varepsilon\}}.
\end{displaymath}
 Given $X,Y$, 
let ${\tilde {\mathbb I}}_{1},\dots, {\tilde {\mathbb I}}_{m}$ be conditionally independent
Bernoulli random variables with success probabilities 
\begin{displaymath}
{\tilde \PP}({\tilde {\mathbb I}}_{k}=1)=1-{\tilde \PP}({\tilde {\mathbb I}}_{k}=0)=\gamma_{k}.
\end{displaymath}
We assume that, given $X,Y$, 
the sequences 
$\{{\mathbb I}_k\}_{k=1}^m$, 
$\{{\tilde {\mathbb I}}_{k}\}_{k=1}^m$  and 
$\{\xi_{3 k}\}_{k=1}^m$
are conditionally independent. Introduce random variables
\begin{displaymath}
L_4=\sum_{1\le k\le m}{\mathbb I}_k\xi_{3 k},
\qquad
L_5=\sum_{1\le k\le m}{\mathbb I}_k{\mathbb I}'_{k}\xi_{3 k},
\qquad
L_6=\sum_{1\le k\le m}{\mathbb I}_k{\tilde {\mathbb I}}_{k}.
\end{displaymath}
Furthermore, we define the random variable $L_7$ as follows. We first generate $X,Y$. 
Then, given $X,Y$, we
generate a Poisson random variable with the conditional mean value 
$\gamma$. The realized value of the Poisson random variable is denoted $L_7$. Thus, we have
$\PP(L_7=r)=\E e^{-\gamma}\gamma^r/r!$, for $r=0,1,\dots$.

We note that $L$ and $L_3$ have the same asymptotic distribution (if any), 
by (\ref{DeltaT}), (\ref{L-L}). Now we prove 
that $L_3$ converges in distribution to $\Lambda_3$. For this purpose we show  that
for any $\varepsilon\in (0,1)$
\begin{eqnarray}\label{TL1}
&&
d_{TV}(L_3, L_4)=o(1),
\qquad
\E(L_4-L_5)=o(1),
\\
\label{TL2}
&&
d_{TV}(L_5, L_6)\le a_2b_1^2 \varepsilon,
\qquad
d_{TV}(L_6,L_7)=o(1),
\\
\label{TL3}
&&
\E e^{itL_7}-\E e^{it\Lambda_3}=o(1).
\end{eqnarray}

Let us prove (\ref{TL1}), (\ref{TL2}), (\ref{TL3}). The first bound of (\ref{TL1}) 
is obtained in the same way as the first bound of (\ref{DeltaT}). To show the second 
bound  of (\ref{TL1}) we invoke the   inequality
\begin{displaymath}
% \E(L_4-L_5)=
{\tilde \E}(L_4-L_5)
=
\sum_{1\le k\le m}(1-{\mathbb I}_k'){\tilde \E}{\mathbb I}_{k1}{\tilde \E}\xi_{3k}
\le
Y_1b_1m^{-1}\sum_{1\le k\le m}(1-{\mathbb I}_k')X_k^2
\end{displaymath}
and obtain
\begin{displaymath}
\E (L_4-L_5)=\E{\tilde  \E}(L_4-L_5)
\le 
b_1^2\E X_1^2{\mathbb I}_{\{X_1\ge \varepsilon b_1^{-1}\beta_n^{1/2}\}}=o(1). 
\end{displaymath}
We note that the right hand side tends to zero since $\beta_n\to+\infty$.

Let us prove the first inequality of (\ref{TL2}).
Proceeding as in (\ref{dTV4}), (\ref{dTV41}) and using the 
identity ${\tilde {\mathbb I}}_k={\tilde {\mathbb I}}_k{\mathbb I}_k'$  we write
\begin{displaymath}
 {\tilde d}_{TV}(L_5, L_6)
\le 
\sum_{1\le k\le m}
{\mathbb I}_k'
{\tilde \PP}({\mathbb I}_k\not=0)
{\tilde d}_{TV}(\xi_{3k}, {\tilde {\mathbb I}}_k).
\end{displaymath}
Next, we estimate 
${\mathbb I}_k'{\tilde d}_{TV}(\xi_{3k}, {\tilde {\mathbb I}}_k)\le \gamma_k^2$, by
LeCam's inequality (\ref{LeCam}), and invoke the inequality
${\tilde \PP}({\mathbb I}_k\not=0)\le \lambda_{k1}$. 
We obtain
\begin{displaymath}
 {\tilde d}_{TV}(L_5, L_6)
\le 
\sum_{1\le k\le m}
{\mathbb I}_k'\lambda_{k1}\gamma_k^2
\le 
\varepsilon\sum_{1\le k\le m}
{\mathbb I}_k'\lambda_{k1}\gamma_k
\le 
\varepsilon Y_1b_1{\hat a}_2.
\end{displaymath}
Here we estimated $\gamma_k^2\le \varepsilon\gamma_k$. Now the inequalities
$d_{TV}(L_5,L_6)\le \E{\tilde d}_{TV}(L_5,L_6)\le a_2b_1^2\varepsilon$
imply the first relation of (\ref{TL2}).

Let us prove the second relation of (\ref{TL2}).
In view of  
(\ref{fact}) it suffices to show that ${\tilde d}_{TV}(L_6,L_7)=o_P(1)$.
For this purpose we write 
\begin{displaymath}
{\tilde d}_{TV}(L_6,L_7)\le 
{\mathbb I}_{{\cal A}_1}{\tilde d}_{TV}(L_6,L_7)+{\mathbb I}_{{\overline {\cal A}}_1},
\end{displaymath}
 where
${\mathbb I}_{{\overline {\cal A}}_1}=o_P(1)$, see (\ref{d-L22}), and estimate 
using  LeCam's inequality (\ref{LeCam})
\begin{displaymath}
{\mathbb I}_{{\cal A}_1}{\tilde d}_{TV}(L_6,L_7)
\le 
{\mathbb I}_{{\cal A}_1}\sum_{1\le k\le m}{\tilde \PP}^2({\mathbb I}_k{\tilde {\mathbb I}}_k=1){\mathbb I}_k'
\le 
b_1^2Y_1^2m^{-2}{\hat a}_4=o_P(1). 
\end{displaymath}
Here we used the fact that $\E X_1^2<\infty$ implies $m^{-2}{\hat a}_4=o_P(1)$.

Finally, we show (\ref{TL3}). We write
${\tilde \E} e^{itL_7}=e^{\gamma(e^{it}-1)}$ and observe that 
\begin{equation}\label{Y1gamma}
 Y_1b_1a_2-\gamma=o_P(1).
\end{equation}
Furthermore, since for any real $t$ the function $z\to e^{z(e^{it}-1)}$ is bounded and uniformly continuous for $z\ge 0$, 
we conclude that (\ref{Y1gamma}) implies the convergence 
\begin{displaymath}
 \E e^{itL_7}=\E e^{\gamma(e^{it}-1)}\to\E e^{Y_1b_1a_2(e^{it}-1)}=\E e^{it\Lambda_3}.
\end{displaymath}
It remains to prove (\ref{Y1gamma}). 
We write $Y_1b_1a_2-\gamma=Y_1b_1(a_2-{\hat a}_2)+Y_1b_1{\hat a}_2-\gamma$ and 
note that
$a_2-{\hat a}_2=o_P(1)$, by the law of large numbers, and
\begin{displaymath}
 0\le \E (Y_1b_1{\hat a}_2-\gamma)
=
b_1^2\E X_1^2{\mathbb I}_{\{X_1\ge \varepsilon b_1^{-1}\beta_n^{1/2}\}}=o(1).
\end{displaymath}

\end{proof}

\bigskip
 
{\it Acknowledgement}. 
%I am the most grateful to Valentas Kurauskas for his valuable comments and for
%coding and running the simulations.
Research of M. Bloznelis
was supported  by the  Research Council of Lithuania grant MIP-053/2011.  

%Author contributions: M.B. designed research and performed theoretical analysis, V.K.  coded and ran the
%simulations.

%\end{document}

\end{document}